OpenStudy (hlambach):

Find a cubic equation with integral coefficients and roots -1/3 and 2+2i. I know I need to use sum and product of the roots, but that's about it.

OpenStudy (noelgreco):

All imaginary roots come in conjugate pairs, so your unwritten root is 2-2i. Set x to each root and then produce a factor x=2+2i so x-2-2i=0 and (x-2-2i) is one factor. Multiply your three factors out and you'll have the cubic polynomial you seek.

OpenStudy (hlambach):

So: (2+2i-x)(x+1/3)(2-2i-x)=0

OpenStudy (hlambach):

But do you know how to do this by using the sum and product?

OpenStudy (noelgreco):

No, I don't.

OpenStudy (hlambach):

Okay, thanks anyways.